In a recent Super Bowl, a TV network predicted that 24 % of the audience would express an interest in seeing one of its forthcoming television shows. The network ran commercials for these shows during the Super Bowl. The day after the Super Bowl, and Advertising Group sampled 149 people who saw the commercials and found that 37 of them said they would watch one of the television shows.

Suppose you are have the following null and alternative hypotheses for a test you are running:

H0:p=0.24
Ha:p>0.24

Calculate the test statistic, rounded to 3 decimal places
Answer #1

Solution :

This hypothesis test is a right tailed test .  

The null and alternative hypothesis is

H0 : p = 0.24

Ha : p > 0.24

\hat p = x / n = 37 / 149 = 0.2483

1 - P0 = 1 - 0.24 = 0.76

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.2483 - 0.24 / [\sqrt(0.24 (0.76)) / 149] = 0.238

The test statistic is 0.238